OpenStudy (anonymous):

A rectangular block of plastic, measuring 1.96 cm by 1.66 cm by 1.94 cm, floats in the middle of the Pacific Ocean. The volume of sea water displaced by the plastic block is 3.77 cm3. Assuming the density of the sea water is 1.025 g/mL, what is the density of the plastic?

OpenStudy (danjs):

Density = mass / Volume Volume block = length x width x height mass of block = mass of water displaced

OpenStudy (danjs):

Mass of displaced Water = Density Of water x Volume Water

OpenStudy (danjs):

Start by calculating, mass of the displaced water...

OpenStudy (anonymous):

volume of the block is 6.31 cm^3

OpenStudy (danjs):

right, Mass of block = mass of water displaced = Density Water x Volume Water \[\frac{ 1.025 ~g~water }{ 1~mL~water }*\frac{ 3.77 ~cm^3~water }{ 1 }*\frac{ 1~mL~water }{ 1~cm^3~water } = (1.025)*(3.77) ~g ~water\]

OpenStudy (danjs):

Density of Plastic = Mass Plastic / Volume Plastic = [(1.025*3.77) / (6.31)] g/cm^3